OpenStudy (anonymous):

Help me find three solutions for theta!! Has to be between 360 degrees and -360 degrees 1.) sin (theta) = sin (15) 2.) sin (theta) = -sin (60) 3.) sin (theta) = 0.5 4.) sin (theta) = -0.71

OpenStudy (anonymous):

Let me try and understand this better so for the first one you have \[\sin \theta= \sin 15\] and you're just trying to find theta? Nothing else, right?

OpenStudy (anonymous):

Correct.

OpenStudy (anonymous):

well if you multiply sin by sin inverse (which looks like this) \[\sin^{-1} \]

OpenStudy (anonymous):

you can cancel out sin

OpenStudy (anonymous):

so if you multiple sin inverse on both sides you get \[\theta=15\]

OpenStudy (anonymous):

Does that make sense?

OpenStudy (anonymous):

Yeah... So I liked at the unit circle and it seems that theta is basically similar around the whole thing

OpenStudy (anonymous):

OK I see you want THREE solutions for theta that makes more sense ok hold on

OpenStudy (anonymous):

Well clearly the first one is 15 degrees, which can also be stated as pi/12 radians, and since sin is the y value and at 30 degrees it is 1/2 and 15 degrees is half of that, it can also be stated as 1/4

OpenStudy (anonymous):

Is that what you are looking for?

OpenStudy (anonymous):

Yes. That makes sense. Thanks!

OpenStudy (anonymous):

Do you need help with the next few or do you understand now?